OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y < -3x + 3 y < x + 2 (1, -5) (1, 5) (5, 1) (-1, 5)

OpenStudy (anonymous):

Y<-3x +3 (x=-1) Y< -3(-1) +3 = 3 + 3 = 6 Y< x+2 (x=5) Y< 5+2=7

OpenStudy (anonymous):

How's that?

OpenStudy (anonymous):

1,-5. ?

OpenStudy (anonymous):

If it's 1 and -5, then it would be different. Y<-3(1) + 3= -3 +3= 0 Y<0 Y< -3(5) + 3 Y< -15 + 3= -12 Y<-12 Of course, that doesn't make a whole lot of sense.